State dalton's law of fartial pressure. Derive it's mathematical from.
1
Expert's answer
2020-04-13T02:34:11-0400

Dalton's law states that in a mixture of non-reacting gases, the total partialpressure exerted is equal to the sum of the partial pressures of the individual gases.

Or, "P=P_1+P_2+P_3+.....+P_n"

Where "P" is total pressure and "P_1,P_2,...,P_n" are the partial pressure of individual component of gases.

Mathematical derivation:

By moles conservation,

"n=n_1+n_2+...+n_n" ......(1)

And "PV=nRT,"

"P_1V=n_1RT" and similarly for all others.

"n=\frac{PV}{RT};n_1=\frac{P_1V}{RT};...;n_n=\frac{P_nV}{RT}"

Substituting these values in equation (1),

"\frac{PV}{RT}=\frac{P_1V}{RT}+\frac{P_2V}{RT}+......+\frac{P_nV}{RT}"

"\implies P=P_1+P_2+...+P_n"